If a country's labor and capital grow at the same rate, is this likely to have the same impact on the growth rate of output? 

  • Yes. If labor and capital are growing at the same rate, the impact on the growth rate of output is the same. 

  • No. Growth in labor always has a bigger impact than growth in capital. 

  • No. Growth in capital always has a bigger impact than growth in labor. 

  • It depends on whether the capital share of output is larger than the labor share of output.


Reference equation: Real GDP per capita growth rate = Nominal GDP per capita growth rate - Inflation rate - Population growth rate 


This equation is an approximation of the exact rate of growth of GDP per capita, and so it results in some errors when calculating this rate. However, the simplified equation both is easy to use and results in small error terms when inflation, nominal GDP growth, and population growth are low, and so it is a useful approximation. The table below lists a fictional country's nominal GDP real GDP, GDP deflator, and population over two years. 

YearNominal GDPGDP deflatorReal GDP(2015 dollars)Poputation
201551,000,00010051,000,0001,000
2016$1,050,000102$1,029,4121,005


a. Verify that the real GDP values in the above table are accurate. 

Instructions: Round your answers to one decimal place. 

b. This country's real GDP per capita for 2015 is_______  This country's real GDP per capita for 2016 is _______ 

c. The growth rate in this country's real GDP per capita between 2015 and 2016 is _______ 
Answer #1

1. The answer is "It depends on whether the capital share of output is larger than the labour share of output"

a. Real GDP in 2015 = 1000000 (same as Nominal GDP as 2015 is base year)

Real GDP in 2016 = Nominal GDP in 2016 * GDP deflator in 2015 / Deflator in 2016

= 1050000 * 100 / 102 = 1029411.76 or (1029412)

b. Real GDP per capital for 2015 = Real GDP in 2015 / Population

= 1000000 / 1000

= 1000

Real GDP per capita for 2016 = Real GDP in 2016 / Population in 2016

= 1029412 / 1005

= 1024.29 or (1024)

c. Growth rate in Real GDP per capita between 2015 and 2016 = GDP per capita in 2016 / GDP per capita in 2015 - 1 * 100

= 1024 / 1000 - 1 * 100

= 2.4 %